OpenStudy (morgengarner):

Which of the following values is a solution of 3|5 + x| = 12? A. 7 B. 1 C. –17 D. –9

OpenStudy (sshayer):

\[\left| 5+x \right|=\frac{ 12 }{ 3 }=4,5+x=\pm4,\] 5+x=4 x=4-5=-1 5+x=-4 x=-4-5=-9